OpenStudy (anonymous):

can someone show me the steps on how to solve to find the zeros of x4-15x2-16, or x3+4x2-3x-12, or x4+2x3+2x2+2x+1

OpenStudy (anonymous):

(Welcome to OpenStudy)

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

Nah I'm too stupid. Sorry dude.

OpenStudy (anonymous):

But I have a few people who can!

OpenStudy (anonymous):

@karatechopper

OpenStudy (anonymous):

I SUMMON THEE!!!

OpenStudy (anonymous):

@uri @Ashleyisakitty @poopsiedoodle @AccessDenied @saifoo.khan @Preetha @Callisto I SUMMON THEE WITH THE POWER OF TAG!!!

OpenStudy (anonymous):

HELP THIS LOST ONE!!!

OpenStudy (anonymous):

Oooooooooommmmm

OpenStudy (anonymous):

Wth this is supposed to work!

OpenStudy (anonymous):

okay? i just really neeed help

OpenStudy (anonymous):

Aha! It worked! Please help!

OpenStudy (karatechopper):

Next time..Pm them..

OpenStudy (karatechopper):

for welcoming..

OpenStudy (anonymous):

@zsoccer needs help, O mighty one..............

OpenStudy (anonymous):

agreed like now please :)

OpenStudy (anonymous):

We tremble in your presence! Aid the first-timer, O magnificent one! All seeing! Wielder of the Ban Hammer!

OpenStudy (anonymous):

your weird.

OpenStudy (anonymous):

WTF where'd he go? Aw well you can't always trust mods.

OpenStudy (anonymous):

Oh well. Now you got no help!

OpenStudy (anonymous):

Sorry! I tried to summon, but I guess they're busy.

Directrix (directrix):

@zsoccer Zeroes of x^4-15x^2-16 are values of x which render the given expression to have a value of zero. First step: Factor: x^4 - 15x^2 - 16

OpenStudy (stamp):

\[x^4-15x^2-16\]\[u=x^2\]\[(x^2)^2-15(x^2)-16\]\[u^2-15u-16\]\[(u-16)(u+1)\]\[(x^2-16)(x^2+1)\]\[(x+4)(x-4)(x^2+1)\]Set equal to zero\[(x+4)(x-4)(x^2+1)=0\]\[x=-4,\ 4,\ i,\ -i\]

OpenStudy (stamp):

verification of above problem's solution: http://www.wolframalpha.com/input/?i=x^4-15x^2-16%3D0

OpenStudy (stamp):

\[x^3+4x^2-3x-12\]\[x^2(x+4)-3(x+4)\]\[(x^2-3)(x+4)\]\[(x+\sqrt{3})(x-\sqrt{3})(x+4)\] Set equal to zero, obtain x = +/- sqrt(3), x = - 4 verification: http://www.wolframalpha.com/input/?i=x^3%2B4x^2-3x-12%3D0

OpenStudy (stamp):

\[x^4+2x^3+2x^2+2x+1\]\[x^4+2x^3+x^2+x^2+2x+1\]\[x^2(x^2+2x+1)+1(x^2+2x+1)\]\[(x^2+1)(x^2+2x+1)\]\[(x^2+1)(x+1)(x+1)\]\[(x^2+1)(x+1)^2\]Set equal to zero\[(x^2+1)(x+1)^2=0\]\[x=-1,\ +/-i\] verification of solutions: http://www.wolframalpha.com/input/?i=x^4%2B2x^3%2B2x^2%2B2x%2B1%3D0